# Temporal-Guided Spiking Neural Networks for Event-Based Human Action Recognition

### v1

@@ I Introduction
Previous studies on SNNs [21, 30, 39, 16] have demonstrated significant achievements in object recognition and classification. We posit that the integration of SNNs and event-based cameras holds exceptional promise for human action recognition (HAR) [34]. Conventional video-based HAR models often raise privacy concerns [4, 5, 38, 17, 20, 25], making them unsuitable for sensitive environments, such as fall detection in private areas like bathrooms. The collaboration of SNNs and event cameras addresses this concern effectively. Event cameras capture only the dynamics of moving subjects without revealing identifiable features or static backgrounds, thus ensuring privacy preservation in HAR tasks.
Recent research [22, 7, 6, 11, 23, 24] has shown SNNs’ effectiveness for HAR on neuromorphic datasets captured by event cameras. However, their handling of long-range temporal information relies exclusively on the SNNs’ spiking neurons, proving inadequate for comprehensive video-based HAR. Effective processing of long-range temporal information is critical for accurate video-based HAR. In light of this, we introduce two novel frameworks to enhance SNNs’ capabilities in this regard. The first, the temporal segment-based SNN (TS-SNN), implements a temporal segmentation strategy [35] on SNNs, segmenting lengthy action sequences into shorter intervals for more efficient long-term data extraction. The second, referred to 3D convolutional SNN (3D-SNN), replaces 2D spatial elements in SNNs with 3D spatio-temporal components, improving temporal information flow across layers.
To advance research in privacy-conscious HAR, we develop the FallingDetection-CeleX dataset, specifically tailored for privacy-preserving applications. At present, there is a scarcity of real-world, event-based HAR datasets with existing ones [1, 22, 27, 9] primarily focus on standard action recognition scenarios, overlooking privacy-sensitive scenarios prevalent in home environments, such as fall detection. Moreover, the majority of these event-based HAR datasets were recorded using DVS128 and DAVIS346 sensors [19], which offer limited resolutions of $128\times 128$ and $346\times 260$ pixels, respectively. In contrast, our dataset employs the high-resolution CeleX-V event camera [2], a 1-megapixel sensor with $1280\times 800$ pixels, capturing more detailed information compared to lower-resolution alternatives. The FallingDetection-CeleX dataset encompasses 875 recordings from 51 subjects performing 7 different actions, including three types of falls.

### v2

@@ I Introduction
Previous studies on SNNs [20, 26, 35, 15] have demonstrated significant achievements in object recognition and classification. We posit that the integration of SNNs and event-based cameras holds exceptional promise for human action recognition (HAR) [30]. Conventional video-based HAR models often raise privacy concerns [4, 5, 34, 16, 19], making them unsuitable for sensitive environments, such as fall detection in private areas like bathrooms. The collaboration of SNNs and event cameras addresses this concern effectively. Event cameras capture only the dynamics of moving subjects without revealing identifiable features or static backgrounds, thus ensuring privacy preservation in HAR tasks.
Recent research [21, 7, 6, 10] has shown SNNs’ effectiveness for HAR on neuromorphic datasets captured by event cameras. However, their handling of long-range temporal information relies exclusively on the SNNs’ spiking neurons, proving inadequate for comprehensive video-based HAR. Effective processing of long-range temporal information is critical for accurate video-based HAR. In light of this, we introduce two novel frameworks to enhance SNNs’ capabilities in this regard. The first, the temporal segment-based SNN (TS-SNN), implements a temporal segmentation strategy [31] on SNNs, segmenting lengthy action sequences into shorter intervals for more efficient long-term data extraction. The second, referred to 3D convolutional SNN (3D-SNN), replaces 2D spatial elements in SNNs with 3D spatio-temporal components, improving temporal information flow across layers.
To advance research in privacy-conscious HAR, we develop the FallingDetection-CeleX dataset, specifically tailored for privacy-preserving applications. At present, there is a scarcity of real-world, event-based HAR datasets with existing ones [1, 21, 23] primarily focus on standard action recognition scenarios, overlooking privacy-sensitive scenarios prevalent in home environments, such as fall detection. Moreover, the majority of these event-based HAR datasets were recorded using DVS128 and DAVIS346 sensors [18], which offer limited resolutions of $128\times 128$ and $346\times 260$ pixels, respectively. In contrast, our dataset employs the high-resolution CeleX-V event camera [2], a 1-megapixel sensor with $1280\times 800$ pixels, capturing more detailed information compared to lower-resolution alternatives. The FallingDetection-CeleX dataset encompasses 875 recordings from 51 subjects performing 7 different actions, including three types of falls.